The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T^2=pie^2L.
If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing?

Answer by checkley75(3666)

8T^2=3.14^2*10
8T^2=9.8596*10
8T^2=98.596
T^2=98.596/8
T^2=12.3245
T=SQRT12.3245
T=3.51 SECONDS FOR THE FULL SWING OF THE CHAIN.